# Hybrid quantum-classical simulations of semiclassical gravity

### 06 Mar 2026 v1

@@ Abstract
Introduction.– Semiclassical quantum field theories (QFTs) dealing with quantum fields evolving on a dynamical classical background occupy a prominent place in contemporary theoretical physics. This type of theories describe phenomena in high-energy physics, including Schwinger pair production Schwinger (1951); Fedotov et al. (2023) or the chiral magnetic effect Kharzeev (2014); Endrődi (2025); in ultrafast and strong-field physics, such as high-harmonic generation and multi-photon ionization Brabec and Krausz (2000); Ciappina et al. (2017); or high-harmonic spectroscopy in condensed matter as a new tool to probe the electronic band structure, topological properties and many-body correlations Kruchinin et al. (2018); Heide et al. (2024). At cosmological scales, the evolution of quantum fluctuations in the classical background of an expanding spacetime and an inflaton field falls also under this semiclassical framework Birrell and Davies (1982); Parker and Toms (2009); Ford (1997); Mukhanov et al. (1992). In the standard model of cosmology, the semiclassical field equations
@@
These considerations motivate the idea of using quantum computers as quantum simulators Feynman (1982); Cirac and Zoller (2012); Altman et al. (2021) for semiclassical field theories. Quantum devices can efficiently encode highly entangled non-Gaussian states of interacting quantum fields, and evolve them accurately in real time Bauer et al. (2023); Di Meglio et al. (2024). In fact, they have already been considered in the context of semiclassical approaches, prominently within QFTs in curved spacetimes Garay et al. (2000); Barceló et al. (2003); Fedichev and Fischer (2004); Lahav et al. (2010); Weinfurtner et al. (2011); Hu et al. (2019); Kolobov et al. (2021); Viermann et al. (2022); Fulgado-Claudio et al. (2023); Švančara et al. (2024); Fulgado-Claudio et al. (2025). In most studies, however, the classical fields are approximated as a fixed dynamical background, neglecting the important yet complicated backreaction Davies et al. (1976); Brout et al. (1995).
In this work, we present a step forward in this direction by developing a hybrid quantum-classical algorithm Peruzzo et al. (2014); Cerezo et al. (2021); Callison and Chancellor (2022) for simulating nonlinear semiclassical field theories. We benchmark it with a scalar–tensor theory as a proxy of gravitational backreaction beyond Eq. (1). The key idea is to encode the quantum sector of the theory onto a quantum device, which implements the real-time quantum field dynamics and allows for measuring a relevant set of observables. When focusing on quantum computers, this step should combine qubit encoding of the fields with Trotter methods Childs et al. (2021) or quantum signal processing Low and Chuang (2019) to evolve them in time. A classical computer is then used to evolve the background fields by numerically solving the classical equations of motion sourced by expectation values of the measured observables. In this process, some subtleties must be addressed, most notably the renormalization of the expectation values and the convergence to the continuum limit. Our closed hybrid loop captures the essence of nonlinear semiclassical dynamics and can be applied to a broad set of problems, from inflation to modified gravity, or even to semiclassical regimes of other disciplines, changing the specific classical equations of motion and the discretized Hamiltonian QFT accordingly.
@@
$+\int\!\!{\rm d}^{2}x\sqrt{-\tilde{g}}\overline{\Psi}\left({\rm i}\tilde{g}_{\mu\nu}\tilde{\gamma}^{\mu}\tilde{\nabla}^{\nu}-m\right)\Psi+\frac{g_{0}^{2}}{2}(\overline{\Psi}\Psi)^{2}.$ (4)
@@
Figure 2: (a) Representation of the effective potential for the classical scalar field in the absence and presence of backreaction. (b) Solutions of the backreaction equations (5)-(6) for the chameleon model (7), in the absence and presence of backreaction. In the latter case, the exact (continuum) and the algorithm-generated solutions are represented. Parameters are fixed, with respect to a reference energy scale $a_{0}$, to $ma_{0}=0.1$, $M_{\rm Pl}=0.5$ (backreaction) or $M_{\rm Pl}=\infty$ (no backreaction), $\lambda a_{0}^{2}=0.5$, $\xi=1/2$, $T/a_{0}=20$, with initial conditions $\phi_{0}=0.5$ and $\dot{\phi}_{0}/a_{0}=0$. The dashed line shows the real-time evolution of the scalar field obtained by the hybrid algorithm. The algorithm parameters are fixed as $L/a_{0}=4$, $a/a_{0}=0.5$, $N_{S}=8$ and $\tilde{\delta t}=0.1$.
@@
Figure 3: (a)-(b) Comparison between the continuum solution and the algorithm solutions for the time evolution of $\phi$ and $\braket{\hat{\overline{\Psi}}\hat{\Psi}}^{\rm ren}$, respectively, for different values of the lattice spacing $a$. The parameters, in $a_{0}$ units, are fixed as $ma_{0}=0.1$, $M_{\rm Pl}=0.5$, $\lambda a_{0}^{2}=0.5$, $\xi=1/2$, $T/a_{0}=5$, with initial conditions for the scalar field $\phi_{0}=0.1$ and $\dot{\phi}_{0}/a_{0}=0$. The algorithm parameters are $L/a_{0}=4$, $\tilde{\delta t}=0.1$. (c) $\mathrm{L}^{2}$ distance between the algorithm-generated time evolution of the classical scalar field and the continuum solution, with matching IR cutoff. The parameters are the same as in (a)-(b). (d) $\mathrm{L}^{2}$ distance between the algorithm-generated time evolution of the classical scalar field in presence and absence of shot noise as a function of the number of shots $N_{\rm shots}$. The plot points are given by the mean of the $\mathrm{L}^{2}$ distance for $N_{\rm rep}=20$ repetitions, and the error bars by their standard deviation. The parameters are fixed, with respect to a reference energy scale $a_{0}$, to $ma_{0}=0.1$, $M_{\rm Pl}=0.5$, $\lambda a_{0}^{2}=0.5$, $\xi=1/2$, $T/a_{0}=10$, with initial conditions for the scalar field $\phi_{0}=0.5$ and $\dot{\phi}_{0}/a_{0}=0$, and algorithm parameters $L/a_{0}=32$, $a/a_{0}=0.5$, $\tilde{\delta t}=0.05$.

### 26 Mar 2026 v2

@@ Abstract
Introduction.– Semiclassical quantum field theories (QFTs) dealing with quantum fields evolving on a dynamical classical background occupy a prominent place in contemporary theoretical physics. This type of theories describe phenomena in high-energy physics, including Schwinger pair production Schwinger (1951); Fedotov et al. (2023) or the chiral magnetic effect Kharzeev (2014); Endrődi (2025); in ultrafast and strong-field physics, such as high-harmonic generation and multi-photon ionization Brabec and Krausz (2000); Ciappina et al. (2017); or high-harmonic spectroscopy in condensed matter as a new tool to probe the electronic band structure, topological properties and many-body correlations Kruchinin et al. (2018); Heide et al. (2024). At cosmological scales, the evolution of quantum fluctuations in the classical background of an expanding spacetime and an inflaton field also falls under this semiclassical framework Birrell and Davies (1982); Parker and Toms (2009); Ford (1997); Mukhanov et al. (1992). In the standard model of cosmology, the semiclassical field equations
@@
These considerations motivate the idea of using quantum computers as quantum simulators Feynman (1982); Cirac and Zoller (2012); Altman et al. (2021) for semiclassical field theories. Quantum devices can efficiently encode highly entangled non-Gaussian states of interacting quantum fields, and evolve them accurately in real time Bauer et al. (2023); Di Meglio et al. (2024). In fact, they have already been considered in the context of semiclassical approaches, prominently within QFTs in curved spacetimes Garay et al. (2000); Barceló et al. (2003); Fedichev and Fischer (2004); Lahav et al. (2010); Weinfurtner et al. (2011); Hu et al. (2019); Kolobov et al. (2021); Viermann et al. (2022); Fulgado-Claudio et al. (2023); Švančara et al. (2024); Fulgado-Claudio et al. (2025), where the general trend has been to treat the classical background as fixed, neglecting the important yet challenging problem of backreaction Davies et al. (1976); Brout et al. (1995). Notable exceptions have addressed platform-specific backreaction in selected analogue systems, including classical-fluid acoustic systems Balbinot et al. (2005b, a); Patrick et al. (2021) and Bose-Einstein condensates within number-conserving formulations Schützhold et al. (2005); Baak et al. (2022); Pal and Fischer (2024), where backreaction enters beyond the leading-order fixed-background description.
In this work, we present a step forward in this direction by developing a general hybrid quantum-classical algorithm Peruzzo et al. (2014); Cerezo et al. (2021); Callison and Chancellor (2022) for simulating nonlinear semiclassical field theories. We benchmark it with a scalar–tensor theory as a proxy of gravitational backreaction beyond Eq. (1). The key idea is to encode the quantum sector of the theory onto a quantum device, which implements the real-time quantum field dynamics and allows for measuring a relevant set of observables. When focusing on quantum computers, this step should combine qubit encoding of the fields with Trotter methods Childs et al. (2021) or quantum signal processing Low and Chuang (2019) to evolve them in time. A classical computer is then used to evolve the background fields by numerically solving the classical equations of motion sourced by expectation values of the measured observables. In this process, some subtleties must be addressed, most notably the renormalization of the expectation values and the convergence to the continuum limit. Our closed hybrid loop captures the essence of nonlinear semiclassical dynamics and can be applied to a broad set of problems, from inflation to modified gravity, or even to semiclassical regimes of other disciplines, changing the specific classical equations of motion and the discretized Hamiltonian QFT accordingly.
@@
$+\int\!\!{\rm d}^{2}x\sqrt{-\tilde{g}}\left[\overline{\Psi}\left({\rm i}\tilde{g}_{\mu\nu}\tilde{\gamma}^{\mu}\tilde{\nabla}^{\nu}-m\right)\Psi+\frac{g_{0}^{2}}{2}(\overline{\Psi}\Psi)^{2}\right].$ (4)
@@
Figure 2: (a) Representation of the effective potential for the classical scalar field in the absence and presence of backreaction. (b) Solutions of the backreaction equations (5)-(6) for the chameleon model (7), in the absence and presence of backreaction. In the latter case, the exact (continuum) and the algorithm-generated solutions are represented. Parameters are fixed, with respect to a reference energy scale $a_{0}$, to $ma_{0}=0.1$, $M_{\rm Pl}=0.5$ (backreaction) or $M_{\rm Pl}=\infty$ (no backreaction), $\lambda a_{0}^{2}=0.5$, $\xi=0.5$, $T/a_{0}=10$, with initial conditions $\phi_{0}=0.1$ and $\dot{\phi}_{0}/a_{0}=0$. The dashed line shows the real-time evolution of the scalar field obtained by the hybrid algorithm. The algorithm parameters are fixed as $L/a_{0}=4$, $a/a_{0}=0.5$, $N_{S}=8$ and $\tilde{\delta t}=0.1$.
@@
Figure 3: (a)-(b) Comparison between the continuum solution and the algorithm solutions for the time evolution of $\phi$ and $\braket{\hat{\overline{\Psi}}\hat{\Psi}}^{\rm ren}$, respectively, for different values of the lattice spacing $a$. The parameters, in $a_{0}$ units, are fixed as $ma_{0}=0.1$, $M_{\rm Pl}=0.5$, $\lambda a_{0}^{2}=0.5$, $\xi=0.5$, $T/a_{0}=5$, with initial conditions for the scalar field $\phi_{0}=0.1$ and $\dot{\phi}_{0}/a_{0}=0$. The algorithm parameters are $L/a_{0}=4$, $\tilde{\delta t}=0.05$. (c) $\mathrm{L}^{2}$ distance between the algorithm-generated time evolution of the classical scalar field and the continuum solution, with matching IR cutoff. The parameters are the same as in (a)-(b). (d) $\mathrm{L}^{2}$ distance between the algorithm-generated time evolution of the classical scalar field in presence and absence of shot noise as a function of the number of shots $N_{\rm shots}$. The plot points are given by the mean of the $\mathrm{L}^{2}$ distance for $N_{\rm rep}=20$ repetitions, and the error bars by their standard deviation. The parameters are fixed, with respect to a reference energy scale $a_{0}$, to $ma_{0}=0.1$, $M_{\rm Pl}=0.5$, $\lambda a_{0}^{2}=0.5$, $\xi=0.5$, $T/a_{0}=10$, with initial conditions for the scalar field $\phi_{0}=0.5$ and $\dot{\phi}_{0}/a_{0}=0$, and algorithm parameters $L/a_{0}=32$, $a/a_{0}=0.5$, $\tilde{\delta t}=0.05$.